In a market, demand is given by P = 100 − Q and the (private) marginal cost of production for the aggregation of all firms (the industry supply curve) is given by MC = Q. Pollution by the industry creates external damages given by the (constant) marginal external cost curve MEC = 30.

(a) Calculate the output and price of if the industry operates under competitive conditions without regulation.

(b) Calculate the socially efficient price and output of the industry

(c) Calculate the deadweight loss due to operating at the competitive level of output rather than the socially efficient level of output.
Answer #1

marker Demand : P = 100- TB - 18 = 1008 gr. MB = d (TB) dg Le 100 - 29 me - 8 which is mpe (marginal private eest) mse marginWeek 8 priee 10/3 Jecantity 160/3 prise are and 100/3 output & 100/3 under competition respectively. by moo mse = me tmee - 1Socially efficient price 10% and e output 55/3 e DWL due to operating at competitive level of output rather than socially eff